OpenStudy (anonymous):

In a lottery game run by a certain state, for every game ticket purchased, a player can pick any 4 numbers from 1 through 43. If the numbers match those drawn by the state, the player wins. If Martin buys 3 tickets, what is his probability of winning?

OpenStudy (anonymous):

Answer Choices: 4/185,115 1/52,890 3/123,410 1/123,410

OpenStudy (ranga):

How many 4 digit numbers can be formed out of 43 numbers? Total Possibilities = 43C4 = 43! / (39! * 4!) = ? If 3 tickets were purchased then the probability of winning is 3 / Total Possibilities

OpenStudy (anonymous):

Wow. Who would have thought that my Honors Algebra 2 teacher would mix the next unit in with the normal distribution unit? Could you tell me the standard form of that formula? Also how you created that 43 C 4 and if there is another thing like that and similar.

OpenStudy (ranga):

If there are n distinct objects and you pick r objects out of it then the total number of ways you can do it is nCr (C is Combinations)\[\Large ^{n}C _{r} = \frac{ n! }{ (n-r)! r! }\]

OpenStudy (anonymous):

So in this case the answer is C. Wow, that seems a bit complex, yet easy.

OpenStudy (ranga):

In the pick-4 lottery, 4 numbers are picked from a total of 43 numbers. n = 43, r = 4 Plug that into the formula to find the total number of ways to pick 4 digit numbers from 43 different numbers. Yeah, C.

OpenStudy (ranga):

Total possibilities = 43! / (39! * 4!) = 43 x 42 x 41 x 40 / (4 x 3 x 2 x 1) = 123410 If 3 tickets are bought the probability of winning is = 3 / 123410

OpenStudy (anonymous):

Question, when you are writing it that way, you don't seem to be including the (n-r)!--r!--that part. Are you just doing the part in the parenthesis? @ranga